# Bridging Information Gaps with Comprehensive Answers: Improving the Diversity and Informativeness of Follow-Up Questions

### v1

@@ 5 Results
To thoroughly assess the quality of the generated follow-up questions, we employ both automatic evaluation (§5.1) and human evaluation (§5.2). As a first step for both evaluations, we automatically identify and remove ungrammatical questions based on syntactic parsing (see Appendix E for a complete description of the filtering process). Table 1 shows the percent of ungrammatical questions that were filtered out for each model. AUG (3.58%) and FULL (6.31%) produce far fewer ungrammatical FQs compared to ORG, demonstrating their ability to generate more well-formed outputs. We focus the rest of our evaluation on the grammatical questions retained after the filtering.
@@ 5.1 Automatic Evaluation
Table 2: Automatic evaluation of follow-up question generation without human reference. AUG and FULL achieve better diversity and readability.
@@
We measure the diversity of the set of follow-up questions generated for a particular (IQ, IA) input, averaged across the dataset. We report Distinct-n Li et al. (2016), which measures the average distinct n-gram across all groups of follow-up questions for a given (IQ, IA) pair. The results in Table 2 show that AUG and FULL perform similarly, and outperform ORG in generating more diverse unigrams and bigrams.
We also apply agglomerative clustering to the sentence embeddings of the FQs (generated with all-mpnet-base-v2) and report the number of clusters formed at a distance threshold of 1.0. We normalize it by the number of generated follow-up questions to quantify the diversity within that set. [^6] The highest score (most diverse) is 1, and the lowest score (least diverse) is achieved for a large set of questions clustered together. Again, Table 2 shows that adding the augmented data substantially improves the diversity of the generated questions.
@@
We report the average question length in terms of the number of tokens. We hypothesize that shorter questions are generally more readable. Table 2 summarizes key statistics: the average length, shortest and longest follow-ups, and standard deviation. The ORG model shows the highest variation in question length (SD = 10.13) compared to FULL (3.77) and AUG (2.98). Notably, its longest follow-up (111 words) far exceeds those from FULL (73) and AUG (24). In contrast, AUG maintains the most consistent length distribution, with a lower standard deviation and a maximum length of 24 words.
Examining the generated follow-up questions, we find that AUG and FULL generally produce concise, well-formed queries, while ORG sometimes generates unclear or uninformative short questions (e.g., “So it’s cultural?”). Meanwhile, the longer questions from ORG and FULL often include extraneous conversational elements, deviating from standard follow-up patterns. Overall, AUG maintains structured, concise outputs, whereas FULL and ORG exhibit greater variability, occasionally producing overly long or conversational responses. More examples are provided in Section 6.1 and Appendix G.
@@
Finally, we use BLEU 1–4 (Papineni et al., 2002), METEOR (Lavie and Agarwal, 2007), and ROUGE-L (Lin, 2004) to measure lexical overlap, and BERTScore (Zhang* et al., 2020) and embedding-based similarity based on the all-mpnet-base-v2 Sentence Transformers Reimers and Gurevych (2019) to measure semantic similarity to the references. In all cases, we compute the maximum score between the human reference and all the generated follow-ups, and report the average for the entire dataset. The automatic evaluation results are summarized in Table 3, revealing a consistent preference to ORG across models. This is unsurprising, given that the questions in the ORG training set and the test set come from the same distribution (the original FollowupQG). The low BLEU scores from the other models align with prior findings in open-ended question generation, where models generate plausible yet lexically diverse outputs that standard n-gram-based metrics fail to capture (Pan et al., 2021). Indeed, the gap between FULL and AUG is less pronounced in BERTScore and sentence similarity, which focus on the semantic alignment with human references and de-emphasize the style.
@@
Table 5: Results of the human evaluation for each aspect. Best results are in bold; statistically significant results are marked with an asterisk.
5.2 Human Evaluation
We conducted a human evaluation to assess the quality of generated follow-up questions across four key aspects: validity, complexity (the level of reasoning required), relevance, and informativeness (Table 4). We randomly sampled 30 (IQ, IA) pairs from the FollowupQG testing set and verified the generated FQs from all models.
The evaluation was carried out using Cloud Connect. To ensure high-quality annotations, we restricted participation to native English-speaking annotators with a minimum of 1,000 completed annotation tasks and an approval rating exceeding 90%. Annotators interacted with a structured evaluation interface (see Appendix H). Each task presented an original question, its corresponding answer, and a generated follow-up question. Annotators first assessed whether the follow-up question was valid. If deemed invalid, they proceeded directly to the next task. Otherwise, they answered four additional evaluation questions, as detailed in Table 4 (See Appendix F for the complete annotation guidelines). Each task was annotated by 3 annotators, yielding substantial inter-annotator agreement, with an average Cohen’s Kappa of $\kappa=0.77$ McHugh (2012).
@@ 5.2 Human Evaluation
Table 5 presents the mean and variance for each evaluation aspect. Overall, AUG achieved the best results across all categories, with a statistically significant difference from the other models (tested with a one-way ANOVA). Over 90% of the FQs generated by AUG were considered valid, and these questions were judged as relevant, somewhat informative, and minimally to moderately complex. FULL closely follows across aspects, while ORG lags behind. The only aspect on which ORG closely follows FULL is relevance, aligning with the findings of Meng et al. (2023) that current models perform well in maintaining relevance. Overall, the results clearly prefer the questions generated by AUG, which excel in validity, complexity, relevance, and informativeness—key qualities for meaningful follow-up questions.
The comparative results across models reveals key insights into the role of data quality versus quantity in the task of follow-up QG. Notably, AUG, trained on the same number of instances as ORG but using only GPT-4-generated questions, consistently outperform both ORG and FULL across most metrics. This suggests that data quality is more critical than dataset size. Despite having ten times more training data, FULL failed to surpass AUG, likely due to noise in the original dataset diluting learning effectiveness. In contrast, AUG benefited from a curated subset of high-quality, reasoning-heavy examples, leading to greater validity, complexity, relevance, and informativeness. These findings challenge the assumption that larger datasets inherently improve performance, highlighting the importance of targeted augmentation with strict quality control. Future work should explore strategies that balance scale with rigorous data refinement to further optimize question generation.
@@ 6.1 Qualitative Analysis
In Table 6, we compare follow-up questions generated by the ORG, AUG, and FULL models for a given (IQ, IA) pair. The questions produced by the ORG model are often either redundant, such as “What is a heuristic?”, or tangential, like responding with “How do you know if you can ELI5?” to the original responder that mentioned they didn’t know if they could explain it to a 5-year-old (ELI5)—diverging from the target concept of heuristics. While the FULL model generates a variety of relevant questions, excelling in diversity, it sometimes includes tangential ones or redundant phrasing, such as “How do heuristics help in segmenting and segmenting information for specific tasks?”, which can affect clarity. In contrast, the AUG model offers the best balance of informativeness and diversity, producing focused and insightful questions, such as “What are some examples of cognitive strategies that rely on heuristics?” and “How does the concept of a heuristic relate to the process of problem-solving?”. Additional examples can be found in Appendix G.
@@ 6.2 Quantifying Information Gain
In Sec. 5.2 we asked annotators to rate how informative each follow-up question is. Here we propose an alternative, automated method for effectively evaluating informativeness without human annotators—by leveraging the “comprehensive answers” generated from GPT-4 (CA; Sec. 3.2). We consider CA as a proxy for the set of all and only the relevant information that can be discussed in the context of (IQ, IA) pair. Thus, an informative FQ cannot be answered by IA (otherwise, the answer to FQ would add no new information); but should be answerable from CA (otherwise, it may be irrelevant). Motivated by this, we prompt GPT-4 to evaluate the answerability of each model’s generated follow-up questions using each IA and CA.
@@
Table 7 reassess the findings from the human evaluation that AUG produced the most informative follow-up questions (36%), followed by FULL (35%) and ORG (25%).
[…]
@@ 7 Conclusion
In this work, we proposed a novel approach to enhance the diversity and informativeness of generated follow-up questions by directly modeling information gaps through a generated comprehensive answer. Training data was generated using GPT-4 and distilled into a smaller model. Our experiments demonstrate that training the smaller model on the augmented dataset significantly outperforms the baselines both in terms of quality and diversity, showing that this method can be effectively adopted to enhance information-seeking dialogues, reduce ambiguities, and improve the accuracy of LLM responses. Future work can explore methods for further optimizing the diversity of generated questions while reducing their redundancy, as well as applying our method to downstream applications that involve multi-turn dialogues. There is also room for developing more automated evaluation metrics to quantify the quality of generated questions, given the cost of human evaluation and the limitations of current automatic evaluation.
@@
We acknowledge several limitations in our work. First, while our CA-based pipeline is effective in knowledge-driven contexts, its applicability to non-knowledge-based conversations, such as opinion-based questions (e.g., “What would you do in such a scenario?”), remains unclear, as the subjective judgment required in these conversations can be difficult for a generated CA to capture. Additionally, although our pipeline prioritizes informativeness, follow-up questions do not always need to introduce new information Kurkul and Corriveau (2018)—for example, requests for simpler explanations (e.g., “Can you explain this in an easier-to-understand way?”). In the future, we hope to extend this method to support various types of follow-up questions and integrate it into downstream dialogue-based applications.
@@
All annotators involved in the human evaluation for this research were fairly compensated, with payment rates exceeding the local minimum wage to ensure equitable remuneration for their time and effort. Prior to recruiting annotators, ethical approval was obtained from the research ethics board at the authors’ institution, ensuring that the human evaluation process adhered to ethical guidelines and that no harm was caused to any individual involved. Additionally, the FollowupQG dataset is publicly available, and we commit to making our dataset, including the augmented data and generated comprehensive answers, accessible to promote transparency and reproducibility in future work.
@@ Appendix A Problematic and Unrelated Instance Examples
Tables 8 and 9 show examples of a problematic instance and an unrelated instance from FollowupQG training set, respectively.
@@ Appendix B LLM Prompts
Table 10: Comprehensive Answer Generation Prompts: GPT-4 first generates an answer from a single perspective, then iteratively provides non-overlapping answers from different perspectives, which are finally synthesized into a unified response.
@@ Appendix C Augmented Data - Human Annotation Guideline
Table 12: Task description and evaluation questions.
@@ C.3 Relevance Question Guideline
Slightly Related Question Example Related Follow-up Question Example
@@ Appendix D Baseline Reproduce
To establish a baseline, we attempted to reproduce the results of Meng et al. (2023) using the reported parameters, as the original implementation was unavailable. We use BART-large, consisting of 24 layers, 16 attention heads, and a hidden dimension of 1024. The initial learning rate (5e-5) led to training instability, which we mitigated by adjusting it to 2e-5 while keeping other hyperparameters unchanged (batch size: 8, epochs: 10, optimizer: Adam Kinga et al. (2015)). The training was conducted on an NVIDIA Tesla V100 GPU with 32GB memory, taking approximately 6 hours per run. We set the random seed as 42. After multiple runs, our reproduced model yielded similar overall performance but with some variation—certain metrics improved while others slightly declined (see Table 16). This model served as the baseline for subsequent experiments.
@@ Appendix F Model Evaluation - Human Annotation Guideline
Table 18: Task description and evaluation questions for BART follow-up question evaluation.

### v2

@@ 5 Results
To thoroughly assess the quality of the generated follow-up questions, we employ both automatic evaluation (§5.1) and human evaluation (§5.2). As a first step for both evaluations, we automatically identify and remove ungrammatical questions based on syntactic parsing (see Appendix E for a complete description of the filtering process). Table 1 shows the percentage of ungrammatical questions that were filtered out for each model. AUG (3.58%) and FULL (6.31%) produce far fewer ungrammatical FQs compared to ORG (33.25%), demonstrating their ability to generate more well-formed outputs. We focus the rest of our evaluation on the grammatical questions retained after the filtering.
@@ 5.1 Automatic Evaluation
Teacher 77.00 95.07 0.869 11.96 3 25 3.19
CoT 80.65 96.34 0.878 16.36 5 30 4.02
Table 2: Automatic evaluation of follow-up question generation without human reference.
@@
We assess the diversity of each set of FQs at the (IQ, IA) level and average the scores across the dataset. First, we report Distinct-n Li et al. (2016), which measures the average distinct n-gram in the FQs associated with each (IQ, IA) pair. Table 2 shows that AUG and FULL achieve comparable Distinct-1/2 scores, both exceeding ORG. Moreover, the AUG Distinct-1/2 scores are comparable to those of the GPT-4o Teacher baseline and only slightly lower than those of the advanced CoT model.
We also compute a sentence-level diversity score. We embed the FQs using all-mpnet-base-v2 Reimers and Gurevych (2019) and apply agglomerative clustering at a distance threshold of 1.0, normalizing the number of clusters by the number of generated follow-up questions. A score of 1 denotes maximum diversity whereas lower values indicate that questions collapse into the same cluster. Again, Table 2 confirms the trend that our augmentation substantially improves the diversity. Moreover, for both metrics, AUG is statistically indistinguishable from the GPT-4o Teacher and CoT baselines, showing that our pipeline elevates small models to LLM-level diversity on the FollowupQG task.
@@
We report the average question length in terms of the number of tokens. We hypothesize that shorter questions are generally more readable. Table 2 lists the average length, shortest and longest follow-ups, and standard deviation (SD). The ORG model shows the greatest variation in question length (SD = 10.13). Notably, its longest follow-up (111 tokens) far exceeds FULL (73) and AUG (24). In contrast, AUG is the most consistent (SD = 2.98; max = 24), with FULL close behind (SD = 3.77)
Qualitatively examining the generated follow-up questions, we find that AUG and FULL generally produce concise, well-formed queries, while ORG sometimes generates very short, vague prompts (e.g., “So it’s cultural?”). Meanwhile, the longer questions from ORG and FULL often include extraneous conversational filler. Overall, AUG maintains structured, concise outputs for follow-up questions, whereas FULL and ORG exhibit greater variability, occasionally producing overly long or conversational phrasing. More examples are provided in §6.1 and Appendix G.
@@
To compare our results with those obtained by Meng et al. (2023), we perform identical automatic evaluations. We measure lexical overlap with BLEU-1-4 (Papineni et al., 2002), METEOR (Lavie and Agarwal, 2007), and ROUGE-L (Lin, 2004), and semantic similarity with BERTScore (Zhang* et al., 2020) and an embedding-based cosine score computed with all-mpnet-base-v2 (Reimers and Gurevych, 2019), following Meng et al. (2023). For each metric, we compute the highest score across all generated follow-ups with the human reference and report the average for the entire dataset. Table 3 shows a consistent advantage for ORG. This is expected, as both its training data and the test set come from the original FollowupQG distribution. The lower BLEU scores for AUG and FULL reflect a common issue in open-ended QG: lexically diverse yet valid questions are under-rewarded by n-gram metrics (Pan et al., 2021). In contrast, the gap between FULL and AUG is much smaller on BERTScore and embedding similarity, which focus on semantic alignment and discount stylistic differences. Consequently, we turn to human evaluation to capture diversity and nuanced informativeness that automatic metrics may overlook.
@@ 5.2 Human Evaluation
Table 5: Human evaluation scores for each aspect comparing ORG, AUG and FULL. Best results are in bold; only statistically significant results are marked with an asterisk.
Teacher AUG CoT
Mean Variance Mean Variance Mean Variance
Validity 0.9457 0.0514 0.9065 0.0849 0.9647 0.0342
Complexity 1.6227 0.6604 1.4798 0.9441 1.5725 0.7575
Relevance 2.1240 0.7928 2.0935 1.0225 2.4863* 0.5500
Informativeness 1.6744 0.6502 1.4517 1.1297 1.6784 0.8095
Table 6: Human evaluation scores for each aspect comparing (i) AUG with Teacher and (ii) AUG with COT GPT-4o. Best results are in bold; only statistically significant results are marked with an asterisk.
We conducted a human evaluation to assess the quality of generated FQs on four criteria: validity, complexity (the level of reasoning required), relevance, and informativeness (Details listed in Table 4). We randomly sampled 30 (IQ, IA) pairs from the FollowupQG testing set and evaluated the generated FQs from all five models.
Human evaluation was conducted on Cloud Connect. To ensure high-quality annotations, we restricted participation to native English-speaking annotators with a minimum of 1,000 completed annotation tasks and an approval rating exceeding 90%. Annotators interacted with a structured evaluation interface (see Appendix H). Each task presented an initial question, its corresponding answer, and a generated FQ candidate. Annotators first assessed whether the FQ was valid. If deemed invalid, they proceeded directly to the next task. Otherwise, they answered four additional evaluation questions, as detailed in Table 4 (See Appendix F for the complete annotation guidelines). Each task was annotated by 3 annotators, yielding substantial inter-annotator agreement, with an average Cohen’s Kappa of $\kappa=0.77$ Cohen (1960).
@@
Table 5 reports the mean and variance of each evaluation criterion for ORG, AUG, and FULL. Overall, AUG achieved the best results across all criteria, with a statistically significant difference from the other models (tested with a one-way ANOVA). Over 90% of the FQs generated by AUG were considered valid, and these questions were rated as relevant, somewhat informative, and minimally to moderately complex. FULL closely follows across aspects, while ORG lags behind. The only aspect on which ORG closely follows FULL is relevance, aligning with the findings of Meng et al. (2023) that current models perform well in maintaining relevance. In sum, the results clearly prefer the questions generated by AUG, which excel in validity, complexity, relevance, and informativeness—qualities essential for meaningful follow-up questions.
To gauge how effectively our information-gap pipeline distills LLM knowledge into smaller models, we repeated human evaluation on the GPT-4o Teacher and CoT baselines described in §4. Table 6 shows that both LLM baselines post slightly higher mean scores than the distilled AUG model, yet the difference is statistically significant only for the Relevance metric. On validity, complexity, and informativeness, AUG—a BART-large model trained on a random, size-matched subset of our augmented data—remains statistically indistinguishable from the much larger Teacher and CoT models. These results underscore the strength of our pipeline: by contrasting initial answers with LLM-generated comprehensive answers, it simulates the human information-seeking process and produces synthetic follow-up questions rich in diversity and informativeness. Fine-tuning on this augmented data enables small, locally deployable models to reach teacher-level quality at a fraction of the inference cost, thereby making high-performance FollowupQG feasible on resource-constrained hardware.
The comparative results across models reveal key insights into the role of data quality versus quantity in the task of FollowupQG. Notably, AUG, trained on the same number of instances as ORG but consisting solely of GPT-4o-generated, high-quality, reasoning-heavy questions, consistently outperforms both ORG and FULL across most metrics, yielding greater validity, complexity, relevance, and informativeness. This indicates that data quality is more critical than dataset size. Despite consisting of ten times more training data, FULL failed to surpass AUG, likely because remaining lower-quality or low-reasoning examples from the original dataset dilute learning. These findings challenge the assumption that larger datasets automatically improve performance and underscore the value of targeted augmentation with strict quality control. Future work should explore strategies to scale data while maintaining rigorous curation to further advance follow-up question generation.
@@ 6.1 Qualitative Analysis
In Table 7, we compare follow-up questions generated by the ORG, AUG, and FULL models for a given (IQ, IA) pair. Questions from ORG are often redundant—e.g., “What is a heuristic?”—or tangential, such as “How do you know if you can ELI5?” to the original responder that mentioned they didn’t know if they could explain it to a 5-year-old (ELI5), thus drifting away from the target concept of heuristics. While the FULL model yields a wider range of relevant questions and excels in diversity, it occasionally produces tangential or wordy phrasing, for instance, “How do heuristics help in segmenting and segmenting information for specific tasks?”, which hurts clarity. By contrast, AUG strikes the best balance of informativeness and diversity, offering focused, insightful questions like “What are some examples of cognitive strategies that rely on heuristics?” and “How does the concept of a heuristic relate to the process of problem-solving?”. Additional examples can be found in Appendix G.
@@ 6.2 Quantifying Information Gain
In §5.2 we asked annotators to rate the informativeness of each follow-up question. We now introduce an automated alternative that requires no human raters, leveraging the GPT-4o “comprehensive answers” (CA; see definition in §3.2). We treat each CA as a proxy for the full body of information relevant to its (IQ, IA) context. An FQ is informative if it (i) cannot be answered from the IA alone—otherwise it adds no new information—and (ii) can be answered from the CA—otherwise it is likely irrelevant. Guided by this rule, we prompt GPT-4o to judge the answerability of every model-generated FQ against both the IA and the corresponding CA.
@@
Table 8 corroborates the human evaluation of informativeness: AUG produces the largest share of informative questions (36 %), followed by FULL (35 %) and ORG (25 %). Comparing the GPT-4o labels with human-annotated informativeness scores (§5.1) further validates the automatic method: annotators assigned higher mean scores to FQs the model classified as informative (1.29) than to those it did not (1.07). A two-sample t-test ($p=0.0011$) confirms the statistical significance, although the effect size is small (Cohen’s $d=0.215$) Cohen (2013).
@@ 7 Conclusion
In this work, we proposed a novel approach to enhance the diversity and informativeness of follow-up questions by explicitly modeling information gaps via an LLM-generated comprehensive answer. We augmented the original FollowupQG dataset with GPT-4o and distilled this data into a small, locally deployable BART-large model. Experiments show that our pipeline enables the small model to outperform peer small-model baselines and to perform comparably to GPT-4o baseline models in terms of validity, complexity, relevance, and informativeness—all at a fraction of the inference cost. These results suggest that targeted, high-quality augmentation can be more impactful than merely increasing dataset size. They also demonstrate that our method offers a practical approach for improving information-seeking dialogues—by reducing ambiguities and enhancing LLM responses—even on systems with limited computational resources.
Future work could explore ways to increase follow-up-question diversity while reducing redundancy, and to extend the pipeline to downstream tasks involving multi-turn dialogue. We also encourage research on stronger automated metrics for evaluating question quality, given the high cost of human annotation and the limitations of current automatic measures.
@@
We acknowledge several limitations in our work. First, while our CA-based pipeline is effective in knowledge-driven contexts, its applicability to non-knowledge-based conversations, such as opinion-based questions (e.g., “What would you do in such a scenario?”), remains unclear, as the subjective judgment required in these conversations can be difficult for a generated CA to capture. Additionally, although our pipeline prioritizes informativeness, follow-up questions do not always need to introduce new information Kurkul and Corriveau (2018)—for example, requests for simpler explanations (e.g., “Can you explain this in an easier-to-understand way?”).
Our work also calls for several future works and expansions. For example, our pipeline can be tested and evaluated on languages besides English, including low-resource languages. Moreover, given the pragmatic applicability of this pipeline and its focus on resource-efficiency, it would be pertinent to evaluate the compute-cost tradeoffs to help users make informed decisions. Lastly, our pipeline’s performance can be evaluated on different combinations of Teacher-Student models as well. In the future, we hope to extend this method to support various types of follow-up questions and integrate it into downstream dialogue-based applications.
@@
All annotators involved in the human evaluation for this research were fairly compensated, with payment rates exceeding the local minimum wage to ensure equitable remuneration for their time and effort. Prior to recruiting annotators, ethical approval was obtained from the research ethics board at the authors’ institution, ensuring that the human evaluation process adhered to ethical guidelines and that no harm was caused to any individual involved. Additionally, the FollowupQG dataset used in this work is publicly available, and we also released the new data created in this work, including the augmented data and generated comprehensive answers, to promote transparency and reproducibility in future work.
Acknowledgments
This work was funded, in part, by the Vector Institute, Canada CIFAR AI Chairs program, Accelerate Foundation Models Research Program Award from Microsoft, and an NSERC discovery grant.
@@ Appendix A Problematic and Unrelated Instance Examples
Tables 9 and 10 show examples of a problematic instance and an unrelated instance from FollowupQG training set, respectively.
@@ Appendix B LLM Prompts
Table 11: Comprehensive Answer Generation Prompts: GPT-4o first generates an answer from a single perspective, then iteratively provides non-overlapping answers from different perspectives, which are finally synthesized into a unified response.
@@ Appendix C Augmented Data - Human Annotation Guideline
Table 13: Task description and evaluation questions used for human annotation of augmented data.
@@ C.3 Relevance Question Guideline
Slightly Related Question Example Not Related Follow-up Question Example
@@ Appendix D Baseline Reproduce
To establish a baseline, we attempted to reproduce the results of Meng et al. (2023) using the reported parameters, as the original implementation was unavailable. We use BART-large, consisting of 24 layers, 16 attention heads, and a hidden dimension of 1024. The initial learning rate (5e-5) led to training instability, which we mitigated by adjusting it to 2e-5 while keeping other hyperparameters unchanged (batch size: 8, epochs: 10, optimizer: Adam Kinga et al. (2015)). The training was conducted on an NVIDIA Tesla V100 GPU with 32GB of memory, taking approximately 6 hours per run. We set the random seed to 42. After multiple runs, our reproduced model yielded similar overall performance but with some variation-certain metrics improved while others slightly declined (see Table 17). This model served as the baseline for subsequent experiments.
@@ Appendix F Model Evaluation - Human Annotation Guideline
Table 19: Task description and evaluation questions used for BART follow-up question evaluation.